OpenStudy (anonymous):

what is the square root of 81x to the 6th power, y to the 3rd power?

OpenStudy (anonymous):

what is square root of 81??

OpenStudy (anonymous):

\[\sqrt{(81x)^6 \cdot y^3}\]??

OpenStudy (anonymous):

No its Square root of 81X to the 6th Power Y to the 3rd Power?

OpenStudy (anonymous):

\[\sqrt{81 \cdot x^6 \cdot y^3}\]

OpenStudy (anonymous):

Use the equation button below and write what is your question??

OpenStudy (anonymous):

@ Cashmere, that is correct without the multiplication signs in between.

OpenStudy (anonymous):

LOLOL :D

OpenStudy (anonymous):

When you don't write anything between them it means multiplication by default

OpenStudy (anonymous):

xy = x.y

OpenStudy (anonymous):

The multiplication sign is inferred to be there. Let's break it up, by the distributive property. \[\sqrt{81} \sqrt{x^6} \sqrt{y^3}\] You know what the square root of 81 is. Remember that\[\sqrt{x} = x^{1/2}\]Hence, we get\[x^{6/2}\] and \[y^{3/2}\] Putting it all back together, we get\[9x^{6/2} y^{3/2}\] What is 6/2?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Exactly. So when what is the exponent on x?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Exactly. So then our final expression would be?